OpenStudy (anonymous):

if the slope of the line passing through the points (-k, 3) and (5,6) is -2, what is the value of k?

OpenStudy (anonymous):

ge ge g eg eg e ge ge ge ge ge ge g sagot agad

OpenStudy (gorv):

slope = (y2-y1)/(x2-x1)

OpenStudy (gorv):

-2= (y2-y1)/(x2-x1) (x1,y1)=(-k,3) (x2,y2)=(5,6) @Henryrenegado put all the values